# Anomaly Detection in Complex Dynamical Systems: A Systematic Framework Using Embedding Theory and Physics-Inspired Consistency

### v1

@@ III-C3 Temporal Differential Consistency informed Autoencoder
where $\Delta t$ represents a small step size. This method is often preferred over forward or backward differences due to its higher accuracy ($O(\Delta t^{2})$), meaning that the approximation error decreases quadratically as $\Delta t$ decreases. Moreover, evaluating points symmetrically around the target reduces truncation errors and improves numerical stability, making the central difference method a reliable choice for derivative estimation [29].
@@
To accurately capture temporal dynamics, we incorporate the central difference method into the training framework of the latent space. The latent representations at the previous ($t-1$) and next ($t+1$) time steps are used to approximate the first-time derivative of the static latent variables ($z$) using the central difference formula. This derivative is then used as a target for the dynamic latent variables ($\dot{z}$), ensuring consistency with the central difference approximation, scaled by the time interval $\Delta t$. A schematic representation of this approach is shown in Fig. 4.
@@
To incorporate temporal dynamics into the training process, we introduce a temporal differential consistency loss (TDC-Loss). This loss enforces consistency between the approximated time derivative of the static latent variables ($z$) and the output of the corresponding dynamic latent variables ($\dot{z}$). During training, the central difference method is used to estimate the time derivative of $z$, which is then compared with $\dot{z}$. The TDC-Loss is combined with the standard reconstruction loss, ensuring that the autoencoder learns a latent representation that captures both state and derivative information. A compact version of the pseudocode for TDC-informed training is presented in Algo. 1.
2:  for each epoch in training_epochs do
3:     for each batch in training_data do
[…]
10:        Update the Autoencoder parameters using the optimizer
11:     end for
12:  end for
III-C4 Consistency KPIs
@@
We employ an embedding based on state-derivative pairs, where the derivative is approximated using the central difference method. While numerical differentiation is generally ill-posed, particularly in the presence of noise [30], our approach does not seek a fully precise description of the system’s dynamics. Instead, we aim for a causal approximation that reliably indicates when the system deviates from normal variability. By choosing a sufficiently large embedding dimension, we ensure that the system operates on a simple attractor geometry. This allows us to expect a nearly constant trendline for both the state and its derivative. To evaluate this, we introduce two key performance indicators (KPIs).
@@
To leverage the causal approximation between state and derivative, the variation of both quantities must remain in a comparable range. We quantify this using the ratio of their z-score normalized variances:
$\eta=\frac{\sigma_{\dot{x}_{z}}^{2}}{\sigma_{x_{z}}^{2}},$ (7)
where $\sigma^{2}$ represents the variance of the z-score normalized values. The interpretation is as follows:
  • If $\eta\approx 1$, the variations are well-matched, and the central difference approximation is appropriate.
  • If $\eta\ll 1$, the derivative is overly smoothed, possibly suppressing meaningful dynamical features.
  • If $\eta\gg 1$, the derivative is too noisy, necessitating further smoothing or the use of more advanced differentiation techniques.
@@
The second KPI evaluates how well the approximated derivative maintains consistency with the state transitions. Specifically, we compute the mean squared error (MSE) between the integrated derivative and the actual state difference over small time intervals:
@@
III-D Anomaly Detection Logic
We aim to incorporate the developed mathematical foundations into anomaly detection logic. By leveraging state-derivative pairs in the latent space, which approximate the causal underlying process, we assume that normal states can be enclosed within a confined region. Deviations from this region indicate abnormal behavior in the physical system.
Training is conducted using the first 50% of the dataset, while validation is performed on data from 50% to 60%. A threshold for each latent node is determined based on validation data, set as a percentile of the validation split. This percentile value is optimized to maximize a classification performance metric, such as the F1-score. The threshold remains constant across all engines rather than being individually adapted. In testing, only the current time step is used, and the model infers the derivative representation from this single time step. This approach is motivated by the fact that the measurement at time $t$ inherently contains values that correspond to rate changes, effectively carrying this derivative information.
The detection logic follows these principles:
  • If all latent nodes exceed their respective thresholds during normal operation, it is classified as a false positive (FP). Otherwise, it is a true negative (TN).
@@ III-D Anomaly Detection Logic
The detection logic can be adjusted based on specific use cases and risk management strategies, such as requiring multiple nodes to surpass thresholds before classifying an instance as an anomaly.
@@ IV-A Consistency of the Proposed Approach
First, we analyze the consistency of the proposed approach. Fig. 5 presents the evolution of the loss terms over 50 training epochs across five independent attempts. The plot shows both the standard Mean Squared Error (MSE) loss and the newly introduced TDC loss term. We observe consistent convergence in the TDC loss term, indicating stable training behavior.
Furthermore, we conducted a Jacobian rank analysis using Singular Value Decomposition (SVD). Applying a zero-threshold of $\epsilon=10^{-9}$, all samples in the test dataset exhibit a full-rank Jacobian. This result suggests that the mapping $\mathcal{E}$, which projects the measurement space $\mathbb{R}^{k}$ to an embedding space $\mathbb{R}^{n}$, behaves as an immersion, preserving the differential structure.
Next, we evaluate two key performance indicators (KPIs) introduced in this study. The first KPI, denoted as $\sigma$, assesses whether the central difference method provides an appropriate approximation of the first time derivative. This metric was computed individually for each engine in the training split (first 60% of the timesteps). The results indicate values close to 1, with deviations in the order of $10^{-7}$, confirming the validity of the approximation.
The second KPI evaluates the relationship between state derivatives, reflecting how well the causal approximation holds across latent nodes in both normal and abnormal conditions. Table I summarizes the mean $\pm$ standard deviation values of the correlation coefficient $\rho$ in both normal and anomalous conditions.
@@ IV-B Phase Space Representation and State-Derivative Relationships
Figure 6: (a) Phase-space plot of the latent node pair (2,6) average. (b) State-derivative consistency KPI for this pair. (c) A detailed view of (a) (highlighted grey box), showing phase-space trajectories and the attractor region at the beginning of the system dynamics. The timestep color map applies to all figures.
@@
In Fig. 6c, we observe the evolution of the phase space representation over time, transitioning from normal to abnormal behavior. Initially, the phase space exhibits an attractor-like structure around which the dynamics evolve in a stable manner. As the system progresses, we observe increased dispersion and a gradual drift away from the attractor toward the anomalous region. Even while still within the range characterized as normal, the state-derivative approximation remains stable, as seen in Fig. 6a. However, subtle deviations from the attractor’s dynamics can already be detected in the earliest phase of the transition. This aligns with our second hypothesized effect, which suggests that even minor disruptions in the system’s stability could serve as an early indicator of anomalies.
@@ V Conclusion
Our results demonstrate that time correlation can be effectively preserved in a 1D embedding, enabling robust anomaly detection. TDC-AE outperforms benchmarks on the C-MAPSS dataset, surpassing computationally intensive models like LSTMs and Transformers, while achieving a 200x reduction in MAC operations, keeping them below 5,000 MACs. This efficiency makes our approach particularly suitable for lightweight edge computing applications.
Future work will focus on further leveraging the concept of state-derivative pairs by expanding on system-theoretic perspectives and physical modeling principles to make the method accessible for further use cases and other types of dynamical systems.

### v2

@@ III-B3 Temporal Differential Consistency informed Autoencoder
where $\Delta t$ represents a small step size. This method is often preferred over forward or backward differences due to its higher accuracy ($O(\Delta t^{2})$), meaning that the approximation error decreases quadratically as $\Delta t$ decreases [29].
@@
To accurately capture temporal dynamics, we incorporate the central difference method into the training framework of the latent space. The latent representations at the previous ($t-1$) and next ($t+1$) time steps are used to approximate the first-time derivative of the static latent variables ($z$) using the central difference formula. This derivative is then used as a target for the dynamic latent variables ($\dot{z}$), ensuring consistency with the central difference approximation, scaled by the time interval $\Delta t$. A schematic representation of this approach is shown in Fig. 7 in Appendix C.
@@
To incorporate temporal dynamics into the training process, we introduce a temporal differential consistency loss (TDC-Loss). This loss enforces consistency between the approximated time derivative of the static latent variables ($z$) and the output of the corresponding dynamic latent variables ($\dot{z}$). During training, the central difference method is used to estimate the time derivative of $z$, which is then compared with $\dot{z}$. The TDC loss serves as a regularizer for the standard reconstruction loss, ensuring that the autoencoder learns a latent representation that captures both state and derivative information. A compact version of the pseudocode for TDC-informed training is presented in Algo. 1 in Appendix B. We use a highly compact autoencoder with three hidden layers in both the encoder and decoder. For details on implementation and hardware, see Appendix B.
III-B4 Consistency Metrics
@@
We employ an embedding based on state-derivative pairs, where the derivative is approximated using the central difference method. While numerical differentiation is generally ill-posed, particularly in the presence of noise [30], our approach does not seek a fully precise description of the system’s dynamics. Instead, we aim for a causal approximation that reliably indicates when the system deviates from normal variability. By choosing a sufficiently large embedding dimension, we ensure that the system operates on a simple attractor geometry. This allows us to expect a nearly constant trendline for both the state and its derivative. To evaluate this, we introduce two metrics
@@
To ensure a balanced causal approximation between state and derivative, their variations must remain comparable. We define this using the ratio of their min-max scaled variances:
$\eta=\frac{\sigma_{\dot{x}_{\text{scaled}}}^{2}}{\sigma_{x_{\text{scaled}}}^{2}},$ (3)
where $\sigma^{2}$ denotes variance after min-max scaling. If $\eta\approx 1$, the variations are well-matched, making central difference approximation suitable. When $\eta\ll 1$, the derivative is overly smoothed, potentially suppressing meaningful dynamics. Conversely, $\eta\gg 1$ indicates excessive noise in the derivative, requiring additional smoothing or refined differentiation techniques.
@@
The second metric evaluates how well the approximated derivative maintains consistency with the state transitions. Specifically, we compute the mean squared error (MSE) between the integrated derivative and the actual state difference over small time intervals:
@@
III-C Anomaly Detection Logic
We integrate the developed mathematical foundations into anomaly detection by leveraging state-derivative pairs in the latent space. Normal states are assumed to lie within a confined region, with deviations indicating anomalies.
Training uses the first 60% of each engine’s data, with an 80/20 split into training and validation sets. Thresholds for each latent node are set as percentiles of the validation data and optimized using a classification metric (e.g., F1-score). This optimization can be performed based on the classification performance on the validation dataset. The threshold remains fixed across all engines. During testing, the model infers the derivative representation from a single time step, as measurements at time $t$ inherently encode rate changes.
Detection logic:
  • If all latent nodes exceed their thresholds during normal operation, it is classified as a false positive (FP); otherwise, it is a true negative (TN).
@@ III-C Anomaly Detection Logic
Detection rules can be adapted based on specific use cases and risk strategies, such as requiring multiple nodes to exceed thresholds before flagging an anomaly.
@@ IV-A Consistency of the Proposed Approach
First, we examine the consistency of the proposed approach. The loss terms consistently converge over 50 training epochs across five independent runs, indicating stable training behavior. Both the standard Mean Squared Error (MSE) loss and the newly introduced TDC loss term exhibit this convergence. For a visual representation, refer to Fig. 6 in Appendix C. We conducted a Jacobian rank analysis using Singular Value Decomposition (SVD). Applying a zero-threshold of $\epsilon=10^{-9}$, all samples in the test dataset exhibit a full-rank Jacobian. This result suggests that the mapping $\mathcal{E}$, which projects the measurement space $\mathbb{R}^{k}$ to an embedding space $\mathbb{R}^{n}$, behaves as an immersion, preserving the differential structure.
The first metric, $\sigma$, confirmed the validity of the central difference method for approximating the first time derivative, with values ranging from 0.7 to 1.2 across engines in the training split. The second metric assessed the correlation of state derivatives, capturing causal relationships under normal and anomalous conditions. Table I summarizes the mean $\pm$ standard deviation of the correlation coefficient $\rho$ in both cases.
@@ IV-B Phase Space Representation and State-Derivative Relationships
Figure 2: (a) Phase-space plot of the latent node pair (2,6) average. (b) State-derivative consistency metric for this pair. (c) A detailed view of (a) (highlighted grey box), showing phase-space trajectories and the attractor region at the beginning of the system dynamics. The timestep color map applies to all figures.
@@
In Fig. 2c, we observe the evolution of the phase space representation over time, transitioning from normal to abnormal behavior. Initially, the phase space exhibits an attractor-like structure around which the dynamics evolve in a stable manner. As the system progresses, we observe increased dispersion and a gradual drift away from the attractor toward the anomalous region. While in the range characterized as normal, the state-derivative approximation remains stable, as seen in Fig. 2a. However, subtle deviations from the attractor’s dynamics can already be detected in the earliest phase of the transition. This aligns with our second hypothesized effect, which suggests that even minor disruptions in the system’s stability could serve as an early indicator of anomalies.
@@ IV-C Comparison with Literature Benchmarks
Furthermore, a variant of the TDC-based architecture has also been applied to the BATADAL dataset, which models cyber-physical water distribution systems. This approach outperformed the challenge-winning method [31], which is grounded in explicit physical modeling. These results underscore the potential of the proposed framework to generalize across structurally different dynamical systems and application domains [32].
@@ V Conclusion
The results demonstrate that our chosen embedding strategy can effectively capture time correlation from a single time step. TDC-AE outperforms benchmarks on the C-MAPSS dataset, surpassing computationally intensive models like LSTMs and Transformers, while achieving a nearly 200x reduction in MAC operations, keeping them below 3,000 MACs. This efficiency makes our approach particularly suitable for lightweight edge computing.
Appendix A Theoretical Background
A-A General Definition of an Oscillatory Dynamical System
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x},t),$ (5)
$\ddot{\theta}+g(\theta,\dot{\theta})+h(\theta,t)=0,$ (6)
Appendix B Implementation details and numerical methods
B-A The Box-Counting Dimension in a dammped driven Pendulum
To illustrate the concept of fractal dimension, we numerically estimated the box-counting dimension for a slice of the phase space of the damped driven pendulum. We considered the first 100–1000 time steps and applied the box-counting method to a subset of the $(\theta,\dot{\omega})$ plane.
Using a range of box sizes $\epsilon$, we counted the number of occupied boxes $N(\epsilon)$ that contained at least one trajectory point. The chosen values of $\epsilon$ were logarithmically spaced as:
$\epsilon\in\{10^{-2},10^{-1.9},10^{-1.8},\dots,10^{-0.4}\}.$
These values were selected to balance resolution constraints and ensure a meaningful estimation within the given simulation timestep.
The relationship between $\log N(\epsilon)$ and $\log(1/\epsilon)$ was then analyzed through a linear fit, where the slope provides an estimate of the box-counting dimension.
Fig. 5a illustrates the fitting result, and Fig. 5b shows the selected phase-space slice used for the computation.
Figure 5: (a) Log-log plot of the box-counting method, showing the linear fit used to estimate the fractal dimension. (b) Phase-space slice of the damped driven pendulum used for the computation. The schematic shows 48 grid cells with $\epsilon\approx 0.45$. The roughly corresponding value pair is highlighted as a green square in panel (a).
B-B TDC-AE
2:  for each epoch in training_epochs do
3:     for each batch in training_data do
[…]
10:        Update the Autoencoder parameters using the optimizer
11:     end for
12:  end for
B-C Implementation of the TDC-AE and Hardware
TABLE III: Hyperparameters used in training the autoencoder for the C-MAPSS (FD001) dataset.
Hyperparameter Value
Optimizer Adamax
Activation Function Tanh
Number of Dense Layers 6
Architecture 24-24-24-8-24-24-24
Learning Rate 0.003
Batch Size 32
$\alpha$ 100
The experiments were conducted on a system running Windows 11, powered by a 13th Gen Intel Core i7-13700H processor (2.4 GHz, 14 cores, 20 threads), along with 32 GB of RAM and an NVIDIA GeForce RTX 4070 Laptop GPU.
The model was implemented using PyTorch 2.3.0 with CUDA 11.8, leveraging GPU acceleration to enhance computational efficiency and reduce training time. The autoencoder’s architecture and training setup are detailed in Table III.
Appendix C Figures